# Truthfulness of Calibration Measures

### v1

@@ 6.1 A Weaker Lower Bound
$b_{0}\coloneqq 0,\ b_{j}\coloneqq\min\left\{\infty\right\}\cup\left\{t\in[b_{j-1}+1,T]\mid\mathrm{Var}_{b_{j\!-\!1}:t}\geq 1\right\},\ \forall j\geq 1.$
@@
$A_{0}\coloneqq 0,\quad A_{j}-A_{j-1}\coloneqq\mathbbm{1}\left[N_{B_{j}}\geq 1\right]-c\cdot\mathbbm{1}\left[b_{j}<\infty\right]\ (j\geq 1),$
@@ 6.2 The Stronger Lower Bound
$b_{0}\coloneqq 0,\ b_{j}\coloneqq\min\left\{\infty\right\}\cup\left\{t\in[b_{j-1}+1,T]\mid\mathrm{Var}_{b_{j\!-\!1}:t}\geq 1\right\},\ \forall j\geq 1.$
@@ 7 $\mathsf{SSCE}$ is Sound, Complete, and Approximately Truthful
$\mathrm{Var}_{T}\coloneqq\sum_{t=1}^{T}p^{\star}_{t}(1-p^{\star}_{t}),$
over $x\sim\mathcal{D}$, where $p^{\star}_{t}\coloneqq\Pr_{x^{\prime}\sim\mathcal{D}}\left[x^{\prime}_{t}=1\mid x^{\prime}_{1:(t-1)}=x_{1:(t-1)}\right]$. By Theorems 5.7 and 6.5, the truthful forecaster gives
@@ 7.1 A Deterministic Algorithm for $O(\sqrt{T})$ $\mathsf{SSCE}$
Finally, the bound (8) holds by taking an integral over $\varepsilon>0$: shorthanding $X\coloneqq\sup_{i\in[N]}M_{T}^{(f_{i})}$, we have
@@ Appendix A Taxonomy of Existing Calibration Measures
$\Delta_{\alpha}\coloneqq\sum_{t=1}^{T}(x_{t}-p_{t})\cdot\mathbbm{1}\left[p_{t}=\alpha\right].$
@@ A.1 Existing Calibration Measures
$\mathsf{ECE}(x,p)\coloneqq\sum_{\alpha\in[0,1]}\left|\sum_{t=1}^{T}(x_{t}-p_{t})\cdot\mathbbm{1}[p_{t}=\alpha]\right|=\sum_{\alpha\in[0,1]}|\Delta_{\alpha}|.$
@@
$\mathsf{smCE}(x,p)\coloneqq\sup_{f\in\mathcal{F}}\sum_{t=1}^{T}f(p_{t})\cdot(x_{t}-p_{t})=\sup_{f\in\mathcal{F}}\sum_{\alpha\in[0,1]}f(\alpha)\cdot\Delta_{\alpha},$
@@
$\mathsf{CalDist}(x,p)\coloneqq\min_{q\in\mathcal{C}(x)}\|p-q\|_{1},$
@@
~$\widetilde{\mathcal{F}}\coloneqq\left\{f:\mathbb{R}\to\mathbb{R}:\|f\|_{2}^{2}+\|f^{\prime}\|_{2}^{2}\leq 1\right\},$
@@
$\mathsf{kCE}^{\mathsf{Lap}}(x,p)\coloneqq\sup_{f\in\widetilde{\mathcal{F}}}\sum_{t=1}^{T}f(p_{t})\cdot(x_{t}-p_{t}).$
@@
$\mathsf{UCal}(x,p)\coloneqq\sup_{S}\left[\sum_{t=1}^{T}S(x_{t},p_{t})-\inf_{\alpha\in[0,1]}\sum_{t=1}^{T}S(x_{t},\alpha)\right],$
@@ A.4 $O(1)$-$\Omega(\sqrt{T})$ Truthfulness Gap of U-Calibration
We first show that the truthful forecaster has an $\Omega(\sqrt{T})$ U-calibration error. Let random variable $X\coloneqq\sum_{t=1}^{T/2}x_{t}$ denote the number of ones among the first $T/2$ random bits. Note that $X$ follows $\mathsf{Binomial}(T/2,1/2)$. Consider the scoring rule defined as:
@@
where $S(\alpha,\beta)\coloneqq\alpha\cdot S(1,\beta)+(1-\alpha)\cdot S(0,\beta)$ is the linear extension of $S$ to $[0,1]^{2}$.
Let $\ell(\alpha)\coloneqq S(\alpha,\alpha)$ denote the uni-variate form of $S$. The following is a standard fact about proper scoring rules (see e.g., [KLST23, Lemma 1 and Corollary 2]).
@@ A.5 Lack of Completeness
Every scoring rule $S:\{0,1\}\times[0,1]\to[0,1]$ induces a calibration measure $\mathsf{CM}^{(S)}(x,p)\coloneqq\sum_{t=1}^{T}S(x_{t},p_{t})$. [^1] Here, we consider scoring rules with co-domain $[0,1]$, since our definition of calibration measures (in Section 2) requires them to be bounded between 00 and $T$ on length-$T$ sequences. When $S$ is proper, it is easy to show that the resulting $\mathsf{CM}^{(S)}$ is perfectly truthful, i.e., $(1,0)$-truthful.
A drawback of such calibration measures is that they all lack completeness. Concretely, consider the squared loss $S(x,p)\coloneqq(x-p)^{2}$. When the outcomes $x_{1},x_{2},\ldots,x_{T}$ are independent and uniformly random bits, the “right” prediction $p_{t}\equiv 1/2$ gives a total penalty of $T/4$, which is only a constant factor away from the maximum possible penalty of $T$. This violates the completeness property in Definition 2.2. In contrast, as shown in Table 2, almost all the other calibration measures would evaluate to $\ll T$ in this case. Such an asymptotic gap better justifies the intuition that $p_{t}\equiv 1/2$ is a much better prediction than, say, $p_{t}\equiv 0$.
@@ B.1 The Upper Bound Part
To handle the general case that $p^{\star}\in[0,1]^{T}$ is arbitrary, we design an alternative sub-routine, the behavior of which depends on whether the sequence $p^{\star}$ is “sufficiently stationary” in some sense. Let $\mu_{\mathrm{first}}\coloneqq\frac{1}{T/2}\sum_{t=1}^{T/2}p^{\star}_{t}$ and $\mu_{\mathrm{second}}\coloneqq\frac{1}{T/2}\sum_{t=T/2+1}^{T}p^{\star}_{t}$ be the averages of the first and the second halves of the sequence, respectively. Let $\mu=(\mu_{\mathrm{first}}+\mu_{\mathrm{second}})/2$ be the overall average.
  • Case 1: $|\mu_{\mathrm{first}}-\mu|>\operatorname{polylog}(T)/\sqrt{T}$. When $\mu_{\mathrm{first}}$ and $\mu$ are far away, we predict $\alpha\coloneqq\frac{\mu_{\mathrm{first}}+\mu}{2}$ at every step. Without loss of generality, suppose that $\mu_{\mathrm{first}}<\mu$, in which case we have
@@
  • Case 2: $|\mu_{\mathrm{first}}-\mu|\leq\operatorname{polylog}(T)/\sqrt{T}$. When $\mu_{\mathrm{first}}$ and $\mu$ are close, we use a strategy that is more similar to the algorithm of [QZ24]. For the first half of the sequence, we predict $\alpha\coloneqq\mu_{\mathrm{first}}$. Let $\Delta_{\mathrm{first}}\coloneqq\sum_{t=1}^{T/2}(x_{t}-\alpha)$ denote the total bias at time $T/2$. Say that $\Delta_{\mathrm{first}}\geq 0$. Then, we will predict $\beta\coloneqq\mu_{\mathrm{second}}+\frac{\Delta_{\mathrm{first}}}{T/2}+\frac{\operatorname{polylog}(T)}{\sqrt{T}}$ in the second half of the sequence. The value of $\beta$ is chosen such that we can offset the bias incurred in the first half (i.e., the $\Delta_{\mathrm{first}}/(T/2)$ term). We also introduce some additional bias (i.e., the $\operatorname{polylog}(T)/\sqrt{T}$ term), so that we can return to a zero bias with high probability. In this case, our sub-routine predicts two different values ($\alpha$ and $\beta$), and they only differ by $\operatorname{polylog}(T)/\sqrt{T}$ with high probability.
@@ 34                   Observe $x_{t}$; $\Delta^{(r)}\leftarrow\Delta^{(r)}+(x_{t}-p_{t})$;
$X_{\mathrm{first}}\coloneqq\sum_{s=t+1}^{t+H}x_{s}\quad\text{and}\quad X\coloneqq\sum_{s=t+1}^{t+2H}x_{s}.$
@@
In the following, we show that a Type 2 round succeeds with probability $1-O(1/T^{(r)})$. Let $X_{\mathrm{first}}\coloneqq\sum_{s=t+1}^{t+H}x_{s}$. Note that $X_{\mathrm{first}}$ is a sum of $H$ independent Bernoulli random variables and $\operatorname*{\mathbb{E}}\left[X_{\mathrm{first}}\right]=\mu_{\mathrm{first}}H$. Furthermore, we have $\Delta_{\mathrm{first}}^{(r)}=X_{\mathrm{first}}-\mu_{\mathrm{first}}H$. By an additive Chernoff bound, we have
@@
Finally, we argue that, with high probability, we exit the second for-loop via the break statement. Let $X_{\mathrm{second}}\coloneqq\sum_{s=t+H+1}^{t+2H}x_{s}$ denote the total outcome in the second half. By symmetry, we only deal with the case that $\Delta_{\mathrm{first}}^{(r)}\geq 0$, where we have $\beta=\mu_{\mathrm{second}}+\Delta_{\mathrm{first}}^{(r)}/H+\sqrt{\frac{\ln T^{(r)}}{2H}}$. If the second for-loop runs all the $H$ iterations in full, at the end of it, the value of $\Delta^{(r)}$ will be given by
@@
Thus, it remains to show that a Type 3 round succeeds with probability $1-O(1/T^{(r)})$. Let $X_{\mathrm{first}}\coloneqq\sum_{s=t+1}^{t+H}x_{s}$. Note that $X_{\mathrm{first}}$ is a sum of independent Bernoulli random variables and $\operatorname*{\mathbb{E}}\left[X_{\mathrm{first}}\right]=\mu_{\mathrm{first}}H$. By a multiplicative Chernoff bound, for any $\delta\geq 0$, we have
@@
Finally, we argue that we exit the second for-loop via the break statement with high probability. Let $X_{\mathrm{second}}\coloneqq\sum_{s=t+H+1}^{t+2H}x_{s}$ denote the total outcome in the second half. Recall that we have $\Delta^{(r)}\geq 0$ at the end of the first for-loop, and that $\beta=\mu_{\mathrm{second}}+\Delta^{(r)}/H+\sqrt{\frac{\ln T^{(r)}}{2H}}$. If the second for-loop runs all the $H$ iterations in full, at the end of it, the value of $\Delta^{(r)}$ will be given by
@@ B.2 The Lower Bound Part
Applying the Berry-Esseen theorem [She10] to the random variable $X\coloneqq\sum_{t=1}^{T}(x_{t}-p^{\star}_{t})$ gives:
@@ C.3 Random Walks with Early Stopping
$X_{t}\coloneqq\sum_{s=1}^{t}\left[\left|x_{s}-p^{\star}_{s}\right|-2p^{\star}_{s}(1-p^{\star}_{s})\right]\cdot\mathbbm{1}\left[p^{\star}_{s}\in\mathcal{I}\right]$
@@ D.3 Auxiliary Lemmas
We have thus proved that the claim also holds for $t-1$. This completes the induction.

### v2

@@ 6.1 A Weaker Lower Bound
$b_{0}\coloneq 0,\ b_{j}\coloneqq\min\left\{\infty\right\}\cup\left\{t\in[b_{j-1}+1,T]\mid\mathrm{Var}_{b_{j\!-\!1}:t}\geq 1\right\},\ \forall j\geq 1.$
@@
$A_{0}\coloneq 0,\quad A_{j}-A_{j-1}\coloneqq\mathbbm{1}\left[N_{B_{j}}\geq 1\right]-c\cdot\mathbbm{1}\left[b_{j}<\infty\right]\ (j\geq 1),$
@@ 6.2 The Stronger Lower Bound
$b_{0}\coloneq 0,\ b_{j}\coloneqq\min\left\{\infty\right\}\cup\left\{t\in[b_{j-1}+1,T]\mid\mathrm{Var}_{b_{j\!-\!1}:t}\geq 1\right\},\ \forall j\geq 1.$
@@ 7 $\mathsf{SSCE}$ is Sound, Complete, and Approximately Truthful
$\mathrm{Var}_{T}\coloneq\sum_{t=1}^{T}p^{\star}_{t}(1-p^{\star}_{t}),$
over $x\sim\mathcal{D}$, where $p^{\star}_{t}\coloneq\Pr_{x^{\prime}\sim\mathcal{D}}\left[x^{\prime}_{t}=1\mid x^{\prime}_{1:(t-1)}=x_{1:(t-1)}\right]$. By Theorems 5.7 and 6.5, the truthful forecaster gives
@@ 7.1 A Deterministic Algorithm for $O(\sqrt{T})$ $\mathsf{SSCE}$
Finally, the bound (8) holds by taking an integral over $\varepsilon>0$: shorthanding $X\coloneq\sup_{i\in[N]}M_{T}^{(f_{i})}$, we have
@@ 8 Discussion
Acknowledgments
This work is supported in part by the National Science Foundation under grants CCF-2145898 and the Graduate Research Fellowship Program under grant DGE 2146752, the Office of Naval Research under grant N00014-24-1-2159, an Alfred P. Sloan fellowship, a Schmidt Sciences AI2050 fellowship, and a Google Research Scholars award. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the funding agencies.
@@ Appendix A Taxonomy of Existing Calibration Measures
$\Delta_{\alpha}\coloneq\sum_{t=1}^{T}(x_{t}-p_{t})\cdot\mathbbm{1}\left[p_{t}=\alpha\right].$
@@ A.1 Existing Calibration Measures
$\mathsf{ECE}(x,p)\coloneq\sum_{\alpha\in[0,1]}\left|\sum_{t=1}^{T}(x_{t}-p_{t})\cdot\mathbbm{1}[p_{t}=\alpha]\right|=\sum_{\alpha\in[0,1]}|\Delta_{\alpha}|.$
@@
$\mathsf{smCE}(x,p)\coloneq\sup_{f\in\mathcal{F}}\sum_{t=1}^{T}f(p_{t})\cdot(x_{t}-p_{t})=\sup_{f\in\mathcal{F}}\sum_{\alpha\in[0,1]}f(\alpha)\cdot\Delta_{\alpha},$
@@
$\mathsf{CalDist}(x,p)\coloneq\min_{q\in\mathcal{C}(x)}\|p-q\|_{1},$
@@
~:-$\widetilde{\mathcal{F}}\coloneq\left\{f:\mathbb{R}\to\mathbb{R}:\|f\|_{2}^{2}+\|f^{\prime}\|_{2}^{2}\leq 1\right\},$
@@
$\mathsf{kCE}^{\mathsf{Lap}}(x,p)\coloneq\sup_{f\in\widetilde{\mathcal{F}}}\sum_{t=1}^{T}f(p_{t})\cdot(x_{t}-p_{t}).$
@@
$\mathsf{UCal}(x,p)\coloneq\sup_{S}\left[\sum_{t=1}^{T}S(x_{t},p_{t})-\inf_{\alpha\in[0,1]}\sum_{t=1}^{T}S(x_{t},\alpha)\right],$
@@ A.4 $O(1)$-$\Omega(\sqrt{T})$ Truthfulness Gap of U-Calibration
We first show that the truthful forecaster has an $\Omega(\sqrt{T})$ U-calibration error. Let random variable $X\coloneq\sum_{t=1}^{T/2}x_{t}$ denote the number of ones among the first $T/2$ random bits. Note that $X$ follows $\mathsf{Binomial}(T/2,1/2)$. Consider the scoring rule defined as:
@@
where $S(\alpha,\beta)\coloneq\alpha\cdot S(1,\beta)+(1-\alpha)\cdot S(0,\beta)$ is the linear extension of $S$ to $[0,1]^{2}$.
Let $\ell(\alpha)\coloneq S(\alpha,\alpha)$ denote the uni-variate form of $S$. The following is a standard fact about proper scoring rules (see e.g., [KLST23, Lemma 1 and Corollary 2]).
@@ A.5 Lack of Completeness
Every scoring rule $S:\{0,1\}\times[0,1]\to[0,1]$ induces a calibration measure $\mathsf{CM}^{(S)}(x,p)\coloneq\sum_{t=1}^{T}S(x_{t},p_{t})$. [^1] Here, we consider scoring rules with co-domain $[0,1]$, since our definition of calibration measures (in Section 2) requires them to be bounded between 00 and $T$ on length-$T$ sequences. When $S$ is proper, it is easy to show that the resulting $\mathsf{CM}^{(S)}$ is perfectly truthful, i.e., $(1,0)$-truthful.
A drawback of such calibration measures is that they all lack completeness. Concretely, consider the squared loss $S(x,p)\coloneq(x-p)^{2}$. When the outcomes $x_{1},x_{2},\ldots,x_{T}$ are independent and uniformly random bits, the “right” prediction $p_{t}\equiv 1/2$ gives a total penalty of $T/4$, which is only a constant factor away from the maximum possible penalty of $T$. This violates the completeness property in Definition 2.2. In contrast, as shown in Table 2, almost all the other calibration measures would evaluate to $\ll T$ in this case. Such an asymptotic gap better justifies the intuition that $p_{t}\equiv 1/2$ is a much better prediction than, say, $p_{t}\equiv 0$.
@@ B.1 The Upper Bound Part
To handle the general case that $p^{\star}\in[0,1]^{T}$ is arbitrary, we design an alternative sub-routine, the behavior of which depends on whether the sequence $p^{\star}$ is “sufficiently stationary” in some sense. Let $\mu_{\mathrm{first}}\coloneq\frac{1}{T/2}\sum_{t=1}^{T/2}p^{\star}_{t}$ and $\mu_{\mathrm{second}}\coloneqq\frac{1}{T/2}\sum_{t=T/2+1}^{T}p^{\star}_{t}$ be the averages of the first and the second halves of the sequence, respectively. Let $\mu=(\mu_{\mathrm{first}}+\mu_{\mathrm{second}})/2$ be the overall average.
  • Case 1: $|\mu_{\mathrm{first}}-\mu|>\operatorname{polylog}(T)/\sqrt{T}$. When $\mu_{\mathrm{first}}$ and $\mu$ are far away, we predict $\alpha\coloneq\frac{\mu_{\mathrm{first}}+\mu}{2}$ at every step. Without loss of generality, suppose that $\mu_{\mathrm{first}}<\mu$, in which case we have
@@
  • Case 2: $|\mu_{\mathrm{first}}-\mu|\leq\operatorname{polylog}(T)/\sqrt{T}$. When $\mu_{\mathrm{first}}$ and $\mu$ are close, we use a strategy that is more similar to the algorithm of [QZ24]. For the first half of the sequence, we predict $\alpha\coloneq\mu_{\mathrm{first}}$. Let $\Delta_{\mathrm{first}}\coloneq\sum_{t=1}^{T/2}(x_{t}-\alpha)$ denote the total bias at time $T/2$. Say that $\Delta_{\mathrm{first}}\geq 0$. Then, we will predict $\beta\coloneq\mu_{\mathrm{second}}+\frac{\Delta_{\mathrm{first}}}{T/2}+\frac{\operatorname{polylog}(T)}{\sqrt{T}}$ in the second half of the sequence. The value of $\beta$ is chosen such that we can offset the bias incurred in the first half (i.e., the $\Delta_{\mathrm{first}}/(T/2)$ term). We also introduce some additional bias (i.e., the $\operatorname{polylog}(T)/\sqrt{T}$ term), so that we can return to a zero bias with high probability. In this case, our sub-routine predicts two different values ($\alpha$ and $\beta$), and they only differ by $\operatorname{polylog}(T)/\sqrt{T}$ with high probability.
@@ 34                   Observe $x_{t}$; $\Delta^{(r)}\leftarrow\Delta^{(r)}+(x_{t}-p_{t})$;
$X_{\mathrm{first}}\coloneq\sum_{s=t+1}^{t+H}x_{s}\quad\text{and}\quad X\coloneq\sum_{s=t+1}^{t+2H}x_{s}.$
@@
In the following, we show that a Type 2 round succeeds with probability $1-O(1/T^{(r)})$. Let $X_{\mathrm{first}}\coloneq\sum_{s=t+1}^{t+H}x_{s}$. Note that $X_{\mathrm{first}}$ is a sum of $H$ independent Bernoulli random variables and $\operatorname*{\mathbb{E}}\left[X_{\mathrm{first}}\right]=\mu_{\mathrm{first}}H$. Furthermore, we have $\Delta_{\mathrm{first}}^{(r)}=X_{\mathrm{first}}-\mu_{\mathrm{first}}H$. By an additive Chernoff bound, we have
@@
Finally, we argue that, with high probability, we exit the second for-loop via the break statement. Let $X_{\mathrm{second}}\coloneq\sum_{s=t+H+1}^{t+2H}x_{s}$ denote the total outcome in the second half. By symmetry, we only deal with the case that $\Delta_{\mathrm{first}}^{(r)}\geq 0$, where we have $\beta=\mu_{\mathrm{second}}+\Delta_{\mathrm{first}}^{(r)}/H+\sqrt{\frac{\ln T^{(r)}}{2H}}$. If the second for-loop runs all the $H$ iterations in full, at the end of it, the value of $\Delta^{(r)}$ will be given by
@@
Thus, it remains to show that a Type 3 round succeeds with probability $1-O(1/T^{(r)})$. Let $X_{\mathrm{first}}\coloneq\sum_{s=t+1}^{t+H}x_{s}$. Note that $X_{\mathrm{first}}$ is a sum of independent Bernoulli random variables and $\operatorname*{\mathbb{E}}\left[X_{\mathrm{first}}\right]=\mu_{\mathrm{first}}H$. By a multiplicative Chernoff bound, for any $\delta\geq 0$, we have
@@
Finally, we argue that we exit the second for-loop via the break statement with high probability. Let $X_{\mathrm{second}}\coloneq\sum_{s=t+H+1}^{t+2H}x_{s}$ denote the total outcome in the second half. Recall that we have $\Delta^{(r)}\geq 0$ at the end of the first for-loop, and that $\beta=\mu_{\mathrm{second}}+\Delta^{(r)}/H+\sqrt{\frac{\ln T^{(r)}}{2H}}$. If the second for-loop runs all the $H$ iterations in full, at the end of it, the value of $\Delta^{(r)}$ will be given by
@@ B.2 The Lower Bound Part
Applying the Berry-Esseen theorem [She10] to the random variable $X\coloneq\sum_{t=1}^{T}(x_{t}-p^{\star}_{t})$ gives:
@@ C.3 Random Walks with Early Stopping
$X_{t}\coloneq\sum_{s=1}^{t}\left[\left|x_{s}-p^{\star}_{s}\right|-2p^{\star}_{s}(1-p^{\star}_{s})\right]\cdot\mathbbm{1}\left[p^{\star}_{s}\in\mathcal{I}\right]$
@@ D.3 Auxiliary Lemmas
We have thus proved that the claim also holds for $t-1$. This completes the induction. ∎